I am having a hard time with this problem. Can someone help me please with the steps on how it is solved as well? Thank you
Solve by substitution method.
8x+6y=-14
-6x+y=49
It's asking for a solution in an ordered pair. I have tried this problem for the past 30 min and can't seem to get the correct pair and I'm not sure what I'm doing wrong??

Answer by checkley77(12844)

8x+6y=-14
-6x+y=49
y=49+6x Replace x in the first equatuon with (49+6x) & solve for y.
8x+6(49+6x)=-14
8x+294+36x=-14
44x=-14-294
44x=-308
x=-308/44
x=-7 ans.
y=49+6*-7
y=49-42
y=7 ans.
Proof:
8*-7+6*7=-14
-56+42=-14
-14=-14